For each DE in this set, classify its order = the highest number of derivatives of the unknown

function, and type: autonomous vs nonautonomous, linear vs nonlinear. If linear, say whether

it is homogeneous or nonhomogeneous, and whether it has constant coecients or nonconstant

coecients.For each DE in this set, classify its order = the
Answer #1

x\left ( t \right )=C1e^{5t}+2C2e^{-t},    y\left ( t \right )=C1e^{5t}-C2e^{-t}

\frac{dx}{dt}=x+4y,        \frac{dy}{dt}= 2x+3y

\frac{dx}{dt}=5C1e^{5t}-2C2e^{-t}= x+4y

x+4y= 5C1e^{5t}-2C2e^{-t} .........\left ( 1 \right )

\frac{dy}{dt}=5C1e^{5t}+C2e^{-t}= 2x+3y

2x+3y= 5C1e^{5t}+C2e^{-t} .............\left ( 2 \right )

substarcting equation 2 from 1 we get2x+3y-x-4y =C2e^{-t}+C2e^{-t}

x-y = 3C2e^{-t} ..................\left ( 3 \right )

substarcting equation 1 from 2 we get

x+4y-2x-3y=-2C2e^{-t}-C2e^{-t}

x+y= -3C2e^{-t} ...........................\left ( 4 \right )

on adding equation 3 and 4 we get

x-y+x+y = 3C2e^{-t}-3C2e^{-t}=0

we have 2x= 0 or x=0

from equation 4 we have x+y= -3C2e^{-t}

           0+y= -3C2e^{-t}

y= -3C2e^{-t}    and x=0

B) x\left ( t=0 \right ) =0,     y\left ( t=0 \right ) =3

x\left ( t=0 \right )=C1e^{0}+2C2e^{0} =0

C1+2C2= 0 ..................\left ( 5 \right )

y\left ( t=0 \right )=C1e^{0}-C2e^{0} =3

C1-C2 =3 .................\left ( 6 \right )

substarcting equation 5 from 6 we get

C1+2C2-C1+C2 = -3

3C2=-3,    C2=-1

from equation 5       C1+2C2=0

                C1-2=0,    C1=2.

therefore the solution is C1=2,    C2=-1